# An Equivalence Result on the Order of Differentiability in Frobenius’ Theorem

### v1

@@ 4 Application
In fact, any normal solution to (1) is locally Lipschitz by definition, and for a locally Lipschitz function, the Clarke derivative can be defined, even if it is non-convex. Hence, the KKT theorem using the Clarke derivatives remains applicable. However, we present a slightly stronger result in which the Clarke derivative is not used. Hence, using the strong properties of convex functions is unnecessary, at least in this context.
@@
$L(x^{*})=\{x\in U|u8x)\leq u(x^{*})\}.$
Because $x^{*}\in L(x^{*})$, $L(x^{*})$ is nonempty. Choose any $x\in L(x^{*})$. Because $u$ is regular, $x-tg(x)\in C_{1}$ for any sufficiently small $t>0$. Therefore, $L(x^{*})$ is the closure of $C_{1}$ and $C_{1}$ is the interior of $L(x^{*})$. In particular, $C_{1}$ is nonempty, and because $u$ is quasi-concave, $C_{1}$ and $L(x^{*})$ are convex. Since $x^{+}\in C_{2}$ and each $h_{i}$ is convex, $C_{2}$ is also nonempty and convex.
@@
Suppose that $U\subset\mathbb{R}^{n}$ is open, $g:U\to\mathbb{R}^{n}\setminus\{0\}$ is locally Lipschitz and satisfies Jacobi’s integrability condition. Then, the following two conditions are equivalent.
@@
Choose an open neighborhood $Z$ of $z^{*}$ such that $z\in Z$ implies $g(z)\cdot p^{*}>0$. Let $X=u^{-1}(u(z^{*}))\cap Z$ and $Y=X\cap(z^{*}+V^{*})$. If $z\in X$, then $p^{*}\notin T_{z}(X)$, and thus $T_{z}(X)+V^{*}=\mathbb{R}^{n}$. Therefore, $X$ is transversal to $z^{*}+V^{*}$, and thus $Y$ is a $1$-dimensional $C^{1}$ manifold and there exists a $C^{1}$ local parametrization $y(s)$ of $Y$ such that $y(0)=z^{*}$. Let $\Phi(s,a)=y(s)+ap^{*}$ and $\Psi(b,c)=z^{*}+bv+cp^{*}$. Then, $\Psi$ is a $C^{\infty}$ diffeomorphism between $\mathbb{R}^{2}$ and $z^{*}+V^{*}$, and $D(\Psi^{-1}\circ\Phi)(0,0)$ is regular. By the inverse function theorem, there exists two open neighborhoods $U_{1},V_{1}$ of $(0,0)$ such that $\Psi^{-1}\circ\Phi:U_{1}\to V_{1}$ is a $C^{1}$ diffeomorphism. Hence, $\Phi=\Psi\circ(\Psi^{-1}\circ\Phi)$ is a $C^{1}$ bijection on $U_{1}$ such that $\Phi^{-1}=(\Psi^{-1}\circ\Phi)^{-1}\circ\Psi^{-1}$ is also $C^{1}$ on $W_{1}\equiv\Phi(U_{1})$. Because $g$ is locally Lipschitz, we can assume that $g$ is $L$-Lipschitz on $W_{1}$. Moreover, by Lemma 1, $u(z+\tau p^{*})$ is increasing in $\tau$ for any $z\in W_{1}$. Without loss of generality, we can assume that $U_{1}$ is convex.
@@
By definition, $x(t)=y(w(t))+a(t)p^{*}$, and thus $z(t)=y(w(t))$. Because $u(y(w(t)))=u(z^{*})$, $u(x(t))\leq u(z^{*})$, and $u(y(s)+\tau p^{*})$ is increasing in $\tau$, we have that there exists $\delta>0$ such that $|t-t^{*}|<\delta$ implies $a(t)\leq 0$. Because $a(t^{*})=0$, $t^{*}$ is a local maximum point of $a(t)$, and thus $\dot{a}(t^{*})=0$. Moreover, by definition of $y(s)$, $\dot{y}(w(t))\cdot g(z(t))=0$, and thus $b(t)\cdot g(z(t))=0$.
@@
$~\frac{|(b(s)-b(t))\cdot(g(z(s))-g(z(t)))|}{s-t}$
@@
$~\frac{b(t)\cdot(g(z(s))-g(z^{*}(t,s)))}{s-t}$
$\leq$ $~L\|b(t)\|\|p^{*}\|\frac{|a(s)-a(t)-\dot{a}(t)(s-t)|}{s-t}\to 0\mbox{ as }s\to t.$
@@
Sometimes, we need a stronger assumption than quasi-convexity. Suppose that $V\subset\mathbb{R}^{n}$ is convex. A function $u:V\to\mathbb{R}$ is said to be strictly quasi-convex if for any $x,y\in V$ and $t\in(0,1)$, if $x\neq y$, then
@@
By the same arguments as in the proof of Theorem 3, we have that $x$ is a solution to this minimization problem. In particular, if $g(x)\cdot y\geq g(x)\cdot x$, then $u(y)\geq u(x)$. Suppose, in addition, that $u$ is strictly quasi-convex. If there is another solution $y\in V$ to the above problem, then for $z=(1-2^{-1})x+2^{-1}y$, $g(x)\cdot z\geq g(x)\cdot x$ and $u(z)<u(x)$, which is a contradiction. Therefore, $x$ is a unique solution to the above problem, and thus if $g(x)\cdot y\geq g(x)\cdot x$ and $y\neq x$, then $u(x)<u(y)$.
@@
Now, suppose that $u$ is not strictly quasi-convex. Then, there exists $x,y\in V$ and $t\in(0,1)$ such that $x\neq y$ and
@@
Suppose that there exists $x,y\in V$ such that $x\neq y$, $g(x)\cdot y\geq g(x)\cdot x$, and $g(y)\cdot x\geq g(y)\cdot y$. By the contrapositive of (i), we have that $g(x)\cdot y=g(x)\cdot x$ and $g(y)\cdot x=g(y)\cdot y$. Let v=yxv=y-x and $x(t)=x+tv$. Choose any $t\in[0,1]$. Because $g(x)\cdot x(t)=g(x)\cdot x$, $u(x(t))\geq u(x)$. In particular, $u(y)\geq u(x)$. By the same arguments, we have that $u(x(t))\geq u(y)$, and in particular, $u(x)\geq u(y)$. Therefore, $u(x)=u(y)$. Because $u$ is quasi-convex, $u(x(t))=u(x)=u(y)$. Therefore, if $X=u^{-1}(u(x))$, then $v\in T_{x(t)}(X)$, and thus $g(x(t))\cdot v=0$ for all $t\in[0,1]$. This implies that
@@ A.1 The Example in the Introduction
However, this function $u$ is $C^{1}$. Suppose that there exists a normal solution $v$ to (1) around $(0,0)$ such that $v$ is $C^{2}$. Without loss of generality, we can assume that the domain $V$ of $v$ is $(-\varepsilon,\varepsilon)^{2}$, where $\varepsilon>0$ is sufficiently small, and $v$ is regular. By Corollary 2, we can assume that $u$ is defined on $V$ and there exists an increasing $C^{1}$ function $\phi$ such that $v=\phi\circ u$. Note that, because $u(0,x_{2})=x_{2}$, $\phi(c)=v(0,c)$. By Proposition 1, there exists a $C^{1}$ positive function $\mu:V\to\mathbb{R}$ such that $\nabla v=\mu g$ on $V$. Therefore, we have that
@@ A.2 An Example of Non-Convex Solutions
Recall the definition of convexity and quasi-convexity. Suppose that $U\subset\mathbb{R}^{n}$ is convex, and consider a function $f:U\to\mathbb{R}$. Then, $f$ is said to be convex if for any $x,y\in U$ and $t\in[0,1]$,
@@
Next, $f$ is said to be quasi-convex if for any $x,y\in U$ and $t\in[0,1]$,
@@
If there is a convex normal solution to (1) defined on some convex set $U$, then it is quasi-convex. If $U=\mathbb{R}^{n}_{++}$ and $g(x)\gg 0$ for any $x\in U$, then by Corollary 2, we conclude that any normal solution to (1) defined on $U$ is quasi-convex. The problem considered in this subsection is the converse of this problem. Suppose that any normal solution to (1) is quasi-convex. Is there a convex normal solution? The answer is negative. The following counterexample is obtained by Arrow and Enthoven [1].
@@
Hence, any level set of $u$ is a straight line. Suppose that $u(x)\geq u(y)$ and u((1t)x+ty)>u(x)u((1-t)x+ty)>u(x) for some $t\in[0,1]$. Then, $0<t<1$. By the intermediate value theorem, there exists $t^{*}\in(t,1]$ such that $u((1-t^{*})x+t^{*}y)=u(x)$. Because the level set $u^{-1}(u(x))$ is a straight line, we have that u((1t)x+ty)=u(x)u((1-t)x+ty)=u(x), which is a contradiction. Therefore, we have that $u$ is quasi-convex. By Corollary 2, any normal solution to (1) is quasi-convex.

### v2

@@ 4 Application
In fact, any normal solution to (1) is locally Lipschitz by definition, and for a locally Lipschitz function, the Clarke derivative can be defined, even if it is non-convex. Hence, the KKT theorem using the Clarke derivatives remains applicable. However, we present a slightly stronger result in which the Clarke derivative is not used. Thus, using the strong properties of convex functions is unnecessary, at least in this context.
@@
$L(x^{*})=\{x\in U|u(x)\leq u(x^{*})\}.$
Because $x^{*}\in L(x^{*})$, $L(x^{*})$ is nonempty. Choose any $x\in L(x^{*})$. Because $u$ is regular, $x-tg(x)\in C_{1}$ for any sufficiently small $t>0$. Therefore, $L(x^{*})$ is the closure of $C_{1}$ and $C_{1}$ is the interior of $L(x^{*})$. In particular, $C_{1}$ is nonempty, and because $u$ is quasi-convex, $C_{1}$ and $L(x^{*})$ are convex. Since $x^{+}\in C_{2}$ and each $h_{i}$ is convex, $C_{2}$ is also nonempty and convex.
@@
Suppose that $U\subset\mathbb{R}^{n}$ is open, $g:U\to\mathbb{R}^{n}\setminus\{0\}$ is locally Lipschitz and satisfies Jacobi’s integrability condition. Then, the following three conditions are equivalent.
@@
Choose an open neighborhood $Z$ of $z^{*}$ such that $z\in Z$ implies $g(z)\cdot p^{*}>0$. Let $X=u^{-1}(u(z^{*}))\cap Z$ and $Y=X\cap(z^{*}+V^{*})$. If $z\in X$, then $p^{*}\notin T_{z}(X)$, and thus $T_{z}(X)+V^{*}=\mathbb{R}^{n}$. Therefore, $X$ is transversal to $z^{*}+V^{*}$, and thus $Y$ is a $1$-dimensional $C^{1}$ manifold and there exists a $C^{1}$ local parametrization $y(s)$ of $Y$ such that $y(0)=z^{*}$. Let $\Phi(s,a)=y(s)+ap^{*}$ and $\Psi(b,c)=z^{*}+bv+cp^{*}$. Then, $\Psi$ is a $C^{\infty}$ diffeomorphism between $\mathbb{R}^{2}$ and $z^{*}+V^{*}$, and $D(\Psi^{-1}\circ\Phi)(0,0)$ is regular. By the inverse function theorem, there exist two open neighborhoods $U_{1},V_{1}$ of $(0,0)$ such that $\Psi^{-1}\circ\Phi:U_{1}\to V_{1}$ is a $C^{1}$ diffeomorphism. Hence, $\Phi=\Psi\circ(\Psi^{-1}\circ\Phi)$ is a $C^{1}$ bijection on $U_{1}$ such that $\Phi^{-1}=(\Psi^{-1}\circ\Phi)^{-1}\circ\Psi^{-1}$ is also $C^{1}$ on $W_{1}\equiv\Phi(U_{1})$. Because $g$ is locally Lipschitz, we can assume that $g$ is $L$-Lipschitz on $W_{1}$. Without loss of generality, we can assume that $U_{1}$ is convex. By Lemma 1, $u(z+\tau p^{*})$ is increasing in $\tau$ for any $z\in W_{1}$.
@@
By definition, $x(t)=y(w(t))+a(t)p^{*}$, and thus $z(t)=y(w(t))$. Because $u(y(w(t)))=u(z^{*})$, $u(x(t))\leq u(z^{*})$, and $u(y(s)+\tau p^{*})$ is increasing in $\tau$, $|t-t^{*}|<\delta$ implies that $a(t)\leq 0$. Because $a(t^{*})=0$, $t^{*}$ is a local maximum point of $a(t)$, and thus $\dot{a}(t^{*})=0$. Moreover, by definition of $y(s)$, $\dot{y}(w(t))\cdot g(z(t))=0$, and thus $b(t)\cdot g(z(t))=0$.
@@
$~\left|\frac{(b(s)-b(t))\cdot(g(z(s))-g(z(t)))}{s-t}\right|$
@@
$~\left|\frac{b(t)\cdot(g(z(s))-g(z^{*}(t,s)))}{s-t}\right|$
$\leq$ $~L\|b(t)\|\|p^{*}\|\frac{|a(s)-a(t)-\dot{a}(t)(s-t)|}{|s-t|}\to 0\mbox{ as }s\to t.$
@@
Sometimes, we need a stronger assumption than quasi-convexity. Suppose that $V\subset\mathbb{R}^{n}$ is convex. A function $u:V\to\mathbb{R}$ is said to be strictly quasi-convex if, for any $x,y\in V$ and $t\in(0,1)$, if $x\neq y$, then
@@
By the same arguments as in the proof of Theorem 3, we have that $x$ is a solution to this minimization problem. In particular, if $g(x)\cdot y\geq g(x)\cdot x$, then $u(y)\geq u(x)$. Suppose, in addition, that $u$ is strictly quasi-convex. If there is another solution $y\in V$ to the above problem, then for $z=(1-2^{-1})x+2^{-1}y$, $g(x)\cdot z\geq g(x)\cdot x$ and $u(z)<u(x)$, which is a contradiction. Therefore, $x$ is the unique solution to the above problem, and thus if $g(x)\cdot y\geq g(x)\cdot x$ and $y\neq x$, then $u(x)<u(y)$.
@@
Now, suppose that $u$ is not strictly quasi-convex. Then, there exist $x,y\in V$ and $t\in(0,1)$ such that $x\neq y$ and
@@
Suppose that there exist $x,y\in V$ such that $x\neq y$, $g(x)\cdot y\geq g(x)\cdot x$, and $g(y)\cdot x\geq g(y)\cdot y$. By the contrapositive of (i), we have that $g(x)\cdot y=g(x)\cdot x$ and $g(y)\cdot x=g(y)\cdot y$. Let v=yxv=y-x and $x(t)=x+tv$. Choose any $t\in[0,1]$. Because $g(x)\cdot x(t)=g(x)\cdot x$, $u(x(t))\geq u(x)$. In particular, $u(y)\geq u(x)$. By the same arguments, we have that $u(x(t))\geq u(y)$, and in particular, $u(x)\geq u(y)$. Therefore, $u(x)=u(y)$. Because $u$ is quasi-convex, $u(x(t))=u(x)=u(y)$. Therefore, if $X=u^{-1}(u(x))$, then $v\in T_{x(t)}(X)$, and thus $g(x(t))\cdot v=0$ for all $t\in[0,1]$. This implies that
@@ A.1 The Example in the Introduction
This function $u$ is not $C^{2}$, but only $C^{1}$. Suppose that there exists a normal solution $v$ to (1) around $(0,0)$ such that $v$ is $C^{2}$. Without loss of generality, we can assume that the domain $V$ of $v$ is $(-\varepsilon,\varepsilon)^{2}$, where $\varepsilon>0$ is sufficiently small, and $v$ is regular. Therefore, we can assume that $u$ is defined on $V$, and by Corollary 2, there exists an increasing $C^{1}$ function $\phi$ such that $v=\phi\circ u$. Note that, because $u(0,x_{2})=x_{2}$, $\phi(c)=v(0,c)$. By Proposition 1, there exists a $C^{1}$ positive function $\mu:V\to\mathbb{R}$ such that $\nabla v=\mu g$ on $V$. Therefore, we have that
@@ A.2 An Example of Non-Convex Solutions
Recall the definition of convexity and quasi-convexity. Suppose that $U\subset\mathbb{R}^{n}$ is convex, and consider a function $f:U\to\mathbb{R}$. Then, $f$ is said to be convex if, for any $x,y\in U$ and $t\in[0,1]$,
@@
Next, $f$ is said to be quasi-convex if, for any $x,y\in U$ and $t\in[0,1]$,
@@
If there is a convex normal solution to (1) defined on some convex set $U$, then it is quasi-convex. If $U=\mathbb{R}^{n}_{++}$ and $g(x)\gg 0$ for any $x\in U$, then by Corollary 2, any normal solution to (1) defined on $U$ is either quasi-convex or quasi-concave. The problem considered in this subsection is the converse of this problem. Suppose that any normal solution to (1) on $U=\mathbb{R}^{n}_{++}$ is either quasi-convex or quasi-concave. Is there a convex normal solution? The answer is negative. The following counterexample was obtained by Arrow and Enthoven [1].
@@
Hence, any level set of $u$ is a straight line. Suppose that $u(x)\geq u(y)$ and u((1t)x+ty)>u(x)u((1-t)x+ty)>u(x) for some $t\in[0,1]$. Then, $0<t<1$. By the intermediate value theorem, there exists $t^{*}\in(t,1]$ such that $u((1-t^{*})x+t^{*}y)=u(x)$. Because the level set $u^{-1}(u(x))$ is a straight line, we have that u((1t)x+ty)=u(x)u((1-t)x+ty)=u(x), which is a contradiction. Therefore, we have that $u$ is quasi-convex. By Corollary 2, any normal solution to (1) is either quasi-convex or quasi-concave.
@@ A.3 Katzner’s Counterexample
Acknowledgments
This paper is supported by Chuo University Grant for Special Research. I am grateful to Toru Maruyama and Hisatoshi Tanaka for their helpful suggestions.